### v1

# Smooth Zone Barrier Lyapunov Functions for Nonlinear Constrained Control Systems

Abstract
This paper introduces the Smooth Zone Barrier Lyapunov Function (s-ZBLF) for output and full-state constrained nonlinear control systems. Unlike traditional BLF methods, where control effort continuously increases as the state moves toward the constraint boundaries, the s-ZBLF method keeps the control effort nearly zero near the origin, with a more aggressive increase as the system approaches the boundary. However, unlike previous works where control effort was zero within a predefined safe region around the origin, the s-ZBLF overcomes the disadvantage of discontinuous control activation by providing a smooth, gradual increase in control effort as the state nears the constraints. This smooth transition improves continuity in the control response and enhances stability by reducing chattering. Additionally, the s-ZBLF provides the advantage of minimal control effort in regions far from the constraints, reducing energy consumption and actuator wear. Two forms of the s-ZBLF—logarithmic-based and rational-based—are presented. Theoretical analysis guarantees that all system states remain within the defined constraints, ensuring boundedness and stability of the closed-loop system. Simulation results validate the effectiveness of the proposed method in handling constrained nonlinear systems.
Index Terms:
Zone Barrier Lyapunov Function, Constrained Control, Nonlinear System, Adaptive Control, Full State Constraints.
I Introduction
In modern control systems, ensuring that state variables remain within predefined bounds is important for maintaining system stability and safety, particularly in safety-critical applications such as autonomous vehicles, robotics, and aerospace systems. These constraints are often required to avoid system failure, maintain safe operation, or ensure compliance with physical or operational limits. To address these constraints, Barrier Lyapunov Functions (BLFs) have emerged as a powerful tool, providing a systematic approach to enforce hard constraints on the system states [1, 2, 3].
To overcome these limitations, the Zone Barrier Lyapunov Function (zBLF) was introduced, offering an improvement over traditional BLFs by introducing the concept of zones [4]. In zBLF, system states are allowed to move freely within predefined safe zones, and control effort is only applied as the states approach the boundary of these zones. This significantly reduces unnecessary control actions, improves energy efficiency, and reduces actuator wear. The zBLF method has shown great promise, particularly in applications where maintaining efficient control is important, as it allows the system to operate without intervention within safe zones.
Despite the advantages of zBLF, there are some inherent limitations. Most notably, zBLF suffers from discontinuous control activation near the boundary. As the system states transition from the free-moving zone to the constrained region, the control effort can switch on abruptly, leading to sharp control actions. This lack of fine-grained control near the boundaries can result in chattering, control instability, or degraded system performance, particularly in applications requiring smooth and continuous control. Additionally, zBLF’s reliance on strict boundaries can make it challenging to ensure a smooth and gradual increase in control effort as the system approaches the constraints.
To address these challenges, we propose a new approach called the Smooth Zone Barrier Lyapunov Function (s-ZBLF). The s-ZBLF retains the energy-saving advantages of zBLF while addressing the issue of discontinuous control activation. Our approach introduces a smooth transition between the safe zone and the constrained region, providing a gradual increase in control effort as the system state moves closer to the boundary. By making the control effort grow progressively, s-ZBLF ensures continuity in the Lyapunov derivative, which leads to smoother and more stable control behavior. Near the origin, where the error is low, the control effort remains minimal, gradually increasing as the system approaches the constraint boundaries, and applying hard constraints only when the boundary is near. This progression ensures that control actions are smoother, leading to more effective constraint enforcement while maintaining system stability.
The key contributions of this paper are as follows:
  • We propose the s-ZBLF, which introduces smooth transitions in control effort as the system approaches constraint boundaries, addressing the discontinuous control issue present in zBLF.
  • s-ZBLF maintains minimal control effort when the system state is far from the constraint boundaries, while gradually increasing the control effort as the state moves closer to the constraints, ensuring smooth and stable transitions.
  • The proposed method is demonstrated through theoretical analysis and simulation results, showing the performance of smooth control, stability, and constraint handling of s-ZBLF.
II Problem Formulation and Preliminaries
System Description: We consider a class of nonlinear systems in strict-feedback form:
$\dot{x}_{i}=f_{i}(x_{1},\dots,x_{i})+g_{i}(x_{1},\dots,x_{i})x_{i+1},\quad i=1,\dots,n-1,$ (1)
$\dot{x}_{n}=f_{n}(x_{1},\dots,x_{n})+g_{n}(x_{1},\dots,x_{n})u,$ (2)
where $x_{i}\in\mathbb{R}$ represents the system states, $u\in\mathbb{R}$ is the control input, and $y=x_{1}$ is the system output. The functions $f_{i}(\cdot)$ and $g_{i}(\cdot)$ are assumed to be known and smooth.
Assumptions: 1. The desired trajectory $x_{d}(t)$ and its derivatives $x_{d}^{(i)}(t)$, $i=1,\dots,n$, are bounded by:
$|x_{d}(t)|<k_{xd},\quad|x_{d}^{(i)}(t)|<k_{xdi},\quad i=1,\dots,n.$ (3)
These bounds ensure that the reference signal remains within the specified constraints for all time.
Lemma 1 (Barbalat’s Lemma): Consider a differentiable function $f(t)$. If the following conditions hold:
  1. 1.
    The time derivative $\dot{f}(t)$ is uniformly continuous.
  2. 2.
    $\lim_{t\to\infty}f(t)$ is finite.
Then, $\lim_{t\to\infty}\dot{f}(t)=0$.
Control Objective: The goal is to design a control law $u(t)$ such that the system tracks the desired trajectory $x_{d}(t)$, while ensuring the state constraints $|x_{i}(t)|<k_{xi}$ are satisfied, with $i=1$ for output constraints and $i=1,\dots,n$ for full-state constraints. Additionally, all closed-loop signals must remain bounded for $t\geq 0$.
Barrier Lyapunov Functions [1] is a scalar function $V(x)$, defined with respect to the system $\dot{x}=f(x)$ on an open region $D$ containing the origin. The function is continuous, positive definite, and has continuous first-order partial derivatives at every point of $D$. Additionally, it holds that $V(x)\to\infty$ as $x$ approaches the boundary of $D$. Along the solution of $\dot{x}=f(x)$, for $x(0)\in D$, it satisfies $V(x(t))\leq b$ for all $t\geq 0$, where $b$ is some positive constant.
There are several types of BLF, such as tangent-type [5], integral-type [6], and logarithmic-type [7]. One example of a standard BLF is given by the logarithmic form $V(x)=\frac{1}{2}\ln\left(\frac{k^{2}}{k^{2}-x^{2}}\right)$, where $k$ represents the constraint boundary.
Lemma 2 [2]: For any positive constants $k_{i}$, $i=1,2,\dots,n$, let $Z:=\{z\in\mathbb{R}^{n}:|z_{i}|<k_{i},i=1,2,\dots,n\}\subset\mathbb{R}^{n}$ be an open set. Consider the system:
$\dot{z}=h(t,z)$
where $h:\mathbb{R}^{+}\times\mathbb{N}\to\mathbb{R}^{n}$ is piecewise continuous in $t$ and locally Lipschitz in $z$, uniformly in $t$, on +×$\mathbb{R}^{+}\times Z$. Define $Z_{i}:=\{z_{i}\in\mathbb{R}:|z_{i}|<k_{i}\}\subset\mathbb{R}$. Suppose there exist positive definite functions $V_{i}:Z_{i}\to\mathbb{R}^{+}$ (for $i=1,2,\dots,n$), which are continuously differentiable on $Z_{i}$ and satisfy
$V_{i}(z_{i})\to\infty\quad\text{as}\quad z_{i}\to\pm k_{i}.$
Let $V(z):=\sum_{i=1}^{n}V_{i}(z_{i})$, and suppose $z_{i}(0)\in Z_{i}$ for all $i=1,2,\dots,n$. If the inequality:
$\dot{V}=\frac{\partial V}{\partial z}h\leq 0$
holds, then $z_{i}(t)\in Z_{i}$ for all $t\in[0,\infty)$.
Zone Barrier Lyapunov Functions [4] is a scalar function $V(x)$, defined with respect to the system $\dot{x}=f(x)$, where the state space is divided into zones. These zones consist of a free-motion zone, where the control effort is zero ($u=0$), and a boundary zone, where the control action increases as the system state approaches the constraint. The ZBLF is continuous, positive definite, and approaches infinity as the system state $x$ reaches the boundary $k$. The function enforces constraints by ensuring that control action is only applied near the boundary, allowing for zero control interventions in the interior zone, leading to reduced energy consumption and actuator wear when the system operates within the free-motion zone.
To the best of the authors’ knowledge, the only ZBLF developed so far is the logarithmic form: $V(x)=\frac{1}{2}\ln\left(\frac{k^{2}e^{-2b}}{k^{2}-x^{2}}\right),$ where $k$ defines the constraint boundary and $b$ is a design parameter controlling the size of the free-motion zone [4, 8].
In this paper, we modify the ZBLF by introducing a smooth transition in the control effort, where the control remains near zero when the system state is far from the constraints and gradually increases as the state approaches the boundaries. This modification eliminates abrupt changes in control activation, providing smoother control transitions and enhancing system stability and performance.
III Control Design with Smooth Zone Barrier Lyapunov Function
III-A Smooth Zone Barrier Lyapunov Function
Smooth Zone Barrier Lyapunov Function is introduced to provide smooth and continuous control activation, ensuring minimal control effort near the origin and progressively increasing control effort as the state nears the boundary. Unlike ZBLF, where control activation may be abrupt, s-ZBLF provides a gradual increase in control effort as the system state approaches the boundary, eliminating discontinuities and enhancing system stability. For a system state $x$ with constraint $|x|<k$, the following two examples of s-ZBLF are considered:
1. $V(x)=\frac{1}{2\beta}\ln\left(\frac{k^{2}}{k^{2}-x^{2}}\right)$
2. $V(x)=\frac{x^{2}}{2(k^{2}-x^{2})(1+\beta x^{2})}$.
In both examples, $\beta>1$ is selected to ensure that the Lyapunov function $V(x)$ exhibits a slower growth rate near the origin, allowing for minimal control effort when the system state is far from the constraint boundaries.
To simplify the notation, the dependence on state and time variables will be omitted from this point forward, whenever it can be done without causing confusion.
III-B Motivation Example: Control Design for a Second Order SISO System
We consider a second-order nonlinear SISO system in the form:
$\dot{x}_{1}=f_{1}(x_{1})+g_{1}(x_{1})x_{2}$ (4)
$\dot{x}_{2}=f_{2}(x_{1},x_{2})+g_{2}(x_{1})u$ (5)
where $u$ is the control input. Let $z_{1}=x_{1}-x_{d}$ and $z_{2}=x_{2}-\alpha_{1}$, where $\alpha_{1}$ is a stabilizing function. The objective is to design an output constrained control law $u$ that respects the state constraint $|x_{1}(t)|<k_{x1}$ while maintaining system stability. We illustrate both forms of the s-ZBLF presented in the previous section with examples.
III-B1 Logarithmic-Based s-ZBLF
The logarithmic Lyapunov function is given by:
$V_{1}=\frac{1}{2\beta}\ln\left(\frac{k^{2}}{k^{2}-z_{1}^{2}}\right)$ (6)
where $k=k_{x1}-k_{xd}$. The time derivative of $V_{1}$ is:
$\dot{V}_{1}=\frac{z_{1}\dot{z}_{1}}{\beta(k^{2}-z_{1}^{2})}$ (7)
Substitute $\dot{z}_{1}$ using the system dynamics (4), $\dot{V}_{1}$ becomes
$\dot{V}_{1}=\frac{z_{1}\left(f_{1}(x_{1})+g_{1}(x_{1})(z_{2}+\alpha_{1})-\dot{y}_{d}\right)}{\beta(k^{2}-z_{1}^{2})}.$ (8)
The stabilizing function $\alpha_{1}$ is designed as:
$\alpha_{1}=\frac{1}{g_{1}}\left(-f_{1}-\beta\kappa_{1}z_{1}\left(k^{2}-z_{1}^{2}\right)+\dot{y_{d}}\right).$ (9)
Substituting (9) into $\dot{V}_{1}$ gives
$\dot{V}_{1}=-\kappa_{1}z_{1}^{2}+\frac{g_{1}z_{1}z_{2}}{\beta(k^{2}-z_{1}^{2})}$ (10)
We define the augmented Lyapunov function $V_{2}$ as:
$V_{2}=V_{1}+\frac{1}{2}z_{2}^{2}.$ (11)
The time derivative of $V_{2}$ is given by
$\dot{V}_{2}=-\kappa_{1}z_{1}^{2}+\frac{g_{1}z_{1}z_{2}}{\beta(k^{2}-z_{1}^{2})}+z_{2}\dot{z}_{2}$ (12)
Substitute $\dot{z}_{2}$ with the system dynamics (5), $\dot{V}_{2}$ becomes:
$\dot{V}_{2}=-\kappa_{1}z_{1}^{2}+\frac{g_{1}z_{1}z_{2}}{\beta(k^{2}-z_{1}^{2})}+z_{2}\left(f_{2}+g_{2}u-\dot{\alpha_{1}}\right).$ (13)
The control law $u$ is designed as:
$u=\frac{1}{g_{2}}\left(-f_{2}+\dot{\alpha_{1}}-\kappa_{2}z_{2}-\frac{g_{1}z_{1}}{\beta(k^{2}-z_{1}^{2})}\right),$ (14)
then, substituting this into (13) gives
$\dot{V}_{2}=-\kappa_{1}z_{1}^{2}-\kappa_{2}z_{2}^{2}.$ (15)
This ensures $\dot{V}_{2}\leq 0$, establishing the stability of the closed-loop system, and the boundedness of $x_{1}$ considering Assumption 1, and Lemma 2.
III-B2 Rational-Based s-ZBLF
For the rational-based s-ZBLF, the Lyapunov function is defined as:
$V_{1}=\frac{z_{1}^{2}}{2(k^{2}-z_{1}^{2})(1+\beta z_{1}^{2})}$ (16)
The time derivative of $V_{1}$ is given by
$\dot{V}_{1}=\frac{(\beta z_{1}^{4}+k^{2})}{(k^{2}-z_{1}^{2})^{2}(1+\beta z_{1}^{2})^{2}}z_{1}\dot{z}_{1}.$ (17)
Define $N=(\beta z_{1}^{4}+k^{2})$ and $D=(k^{2}-z_{1}^{2})^{2}(1+\beta z_{1}^{2})^{2}$. The stabilizing function $\alpha_{1}$ is:
$\alpha_{1}=\frac{1}{g_{1}}\left(-f_{1}-\kappa_{1}\frac{N}{D}z_{1}+\dot{y_{d}}\right).$ (18)
Substituting $\alpha_{1}$ and $\dot{z}_{1}$, $\dot{V}_{1}$ becomes:
$\dot{V}_{1}=-\kappa_{1}z_{1}^{2}+g_{1}\frac{N}{D}z_{1}z_{2}.$ (19)
Define the augmented Lyapunov function $V_{2}$ as
$V_{2}=V_{1}+\frac{1}{2}z_{2}^{2}.$ (20)
The time derivative of $V_{2}$ is given by
$\dot{V}_{2}=-\kappa_{1}z_{1}^{2}+g_{1}\frac{N}{D}z_{1}z_{2}+z_{2}\left(f_{2}+g_{2}u-\dot{\alpha_{1}}\right).$ (21)
The control law $u$ is designed as:
$u=\frac{1}{g_{2}}\left(-f_{2}+\dot{\alpha_{1}}-\kappa_{2}z_{2}-g_{1}\frac{N}{D}z_{1}\right).$ (22)
Substituting (22) into (21) gives
$\dot{V}_{2}=-\kappa_{1}z_{1}^{2}-\kappa_{2}z_{2}^{2}.$ (23)
This ensures $\dot{V}_{2}\leq 0$, establishing the stability of the closed-loop system, and the boundedness of $x_{1}$ considering Assumption 1, and Lemma 2 for the rational-based s-ZBLF.
III-C s-ZBLF for Higher-Order Nonlinear Systems with Output Constraints
Backstepping is used in the control design in this section. Let $z_{1}=x_{1}-x_{d}$ and $z_{i}=x_{i}-\alpha_{i-1}$, for $i=2,\dots,n$. Define the state constraint as $|x_{1}(t)|<k_{x1}$, and let $k_{1}=k_{x1}-k_{xd}$ is the error constraint. The Lyapunov function candidates are defined using the logarithmic-based s-ZBLF as follows:
$V_{1}=\frac{1}{2\beta}\ln\left(\frac{k_{1}^{2}}{k_{1}^{2}-z_{1}^{2}}\right)$ (24)
$V_{i}=V_{i-1}+\frac{1}{2}z_{i}^{2},\quad i=2,\dots,n$ (25)
where $\beta>0$ is a design parameter. The stabilizing functions and control law are designed as:
$\alpha_{1}=\frac{1}{g_{1}}\left(-f_{1}-\beta\kappa_{1}z_{1}\left(k_{1}^{2}-z_{1}^{2}\right)+\dot{y}_{d}\right)$ (26)
$\alpha_{2}=\frac{1}{g_{2}}\left(-f_{2}+\dot{\alpha}_{1}-\kappa_{2}z_{2}-\frac{g_{1}z_{1}}{\beta\left(k_{1}^{2}-z_{1}^{2}\right)}\right)$ (27)
$\alpha_{i}=\frac{1}{g_{i}}\left(-f_{i}+\dot{\alpha}_{i-1}-\kappa_{i}z_{i}-g_{i-1}z_{i-1}\right),\quad i=3,\dots,n$ (28)
$u=\alpha_{n}$ (29)
where $\kappa_{i}>0$ are constants, and $\dot{\alpha}_{i-1}$ is given by:
$\dot{\alpha}_{i-1}=\sum_{j=1}^{i-1}\frac{\partial\alpha_{i-1}}{\partial x_{j}}\left(f_{j}+g_{j}x_{j+1}\right)+\sum_{j=0}^{i-1}\frac{\partial\alpha_{i-1}}{\partial y^{(j)}_{d}}y^{(j+1)}_{d}$ (30)
Then, the closed-loop system dynamics are:
$\dot{z}_{1}=-\beta\kappa_{1}z_{1}\left(k_{1}^{2}-z_{1}^{2}\right)+g_{1}z_{2}$ (31)
$\dot{z}_{2}=-\kappa_{2}z_{2}-\frac{g_{1}z_{1}}{\beta\left(k_{1}^{2}-z_{1}^{2}\right)}+g_{2}z_{3}$ (32)
$\dot{z}_{i}=-\kappa_{i}z_{i}-g_{i-1}z_{i-1}+g_{i}z_{i+1},\quad i=3,\dots,n-1$ (33)
$\dot{z}_{n}=-\kappa_{n}z_{n}-g_{n-1}z_{n-1}$ (34)
Theorem 1: Consider the nonlinear system described by (4), (5) with stabilizing functions (26)-(28), and the control law (29) under the Assumption 1, and assume the initial conditions satisfy $|z_{1}(0)|<k_{1}$. Let the design parameters $\beta>0$ and $\kappa_{i}>0$ be chosen appropriately. Then, the following properties hold:
(i) The signals $z_{i}(t)$, $i=1,\dots,n$, remain in the compact set defined by:
$\mathcal{Z}=\left\{z\in\mathbb{R}^{n}:|z_{1}|\leq D_{z_{1}},\sum_{i=2}^{n}z_{i}^{2}\leq 2V_{n}(0)\right\}$
where:
$D_{z_{1}}=k_{1}\sqrt{1-e^{-2\beta V_{n}(0)}}$
and $V_{n}(0)$ is the initial value of the Lyapunov function.
(ii) The output $x_{1}(t)$ remains within the constraint set $|x_{1}(t)|<k_{x1}$ for all $t\geq 0$.
(iii) All closed-loop signals are bounded.
(iv) The tracking error $z_{1}(t)=x_{1}(t)-x_{d}(t)$ converges to zero asymptotically as $t\to\infty$, i.e., $\lim_{t\to\infty}z_{1}(t)=0$.
Proof:
(i) From the closed-loop system dynamics (31)-(34) and the Lyapunov function defined by (24)-(25), the time derivative of $V_{n}$ can be obtained $\dot{V}_{n}=-\sum_{i=1}^{n}\kappa_{i}z_{i}^{2}$, ensuring that $\dot{V}_{n}\leq 0$, which implies
$V_{n}(t)\leq V_{n}(0)$
for all $t\geq 0$. Consequently, the signals $z_{i}(t)$, $i=1,\dots,n$, remain bounded. Specifically, from $\frac{1}{2\beta}\ln\left(\frac{k_{1}^{2}}{k_{1}^{2}-z_{1}^{2}}\right)\leq V_{n}(0)$, it follows that $|z_{1}(t)|\leq k_{1}\sqrt{1-e^{-2\beta V_{n}(0)}}$, and the remaining states satisfy $\sum_{i=2}^{n}z_{i}^{2}\leq 2V_{n}(0)$. Therefore, the signals $z_{i}(t)$, $i=1,\dots,n$, remain in the compact set $\mathcal{Z}$.
(ii) From the fact that $z_{1}(t)\leq D_{z_{1}}$, we have $|x_{1}(t)|=|z_{1}(t)+x_{d}(t)|\leq D_{z_{1}}+|x_{d}(t)|$. Since $|x_{d}(t)|\leq k_{xd}$, we conclude that:
$|x_{1}(t)|\leq k_{1}+k_{xd}=k_{x1}$
Hence, the output constraint $|x_{1}(t)|<k_{x1}$ is never violated.
(iii) From (i), we know that the error signals $z_{1}(t),\dots,z_{n}(t)$ are bounded. The boundedness of $z_{1}(t)$ and $x_{d}(t)$ implies that the state $x_{1}(t)$ is bounded. Given that $\dot{x}_{d}(t)$ is bounded by Assumption 1, it follows from (26) that the stabilizing function $\alpha_{1}(t)$ is also bounded. Consequently, since $x_{2}=z_{2}+\alpha_{1}$, the state $x_{2}(t)$ is bounded. From (27), $\alpha_{2}(t)$ is a function of bounded terms, implying that $\alpha_{2}(t)$ is also bounded. This ensures that $x_{3}=z_{3}+\alpha_{2}$ is bounded as well. Following this line of argument, for $i=3,\dots,n-1$, the stabilizing function $\alpha_{i}(t)$ depends on bounded signals, as shown in equation (28), ensuring the boundedness of $\alpha_{i}(t)$ and the state $x_{i+1}(t)$.
Since $u=\alpha_{n}$ from equation (28), and $\alpha_{n}(t)$ is bounded, the control input $u(t)$ is also bounded. Therefore, all closed-loop signals are bounded.
(iv) From (i) and (iii), the signals $z_{i}(t)$, $i=1,\dots,n$, are bounded, and consequently, $V_{n}(t)$ remains bounded. Given that $\dot{V}_{n}(t)$ is non-increasing and bounded, it implies that $\ddot{V}_{n}(t)$ is bounded as well. By applying Barbalat’s Lemma, we conclude that $\dot{V}_{n}(t)\to 0$ as $t\to\infty$, which leads to $z_{i}(t)\to 0$, for $i=1,\dots,n$, ensuring the output tracking error $z_{1}(t)$ converges to zero asymptotically.
III-D s-ZBLF for Higher-Order Nonlinear Systems with Full State Constraints
Denote $z_{1}=x_{1}-x_{d}$ and $z_{i}=x_{i}-\alpha_{i-1}$ for $i=2,\dots,n$. Consider the Lyapunov function candidate:
$V=\sum_{i=1}^{n}V_{i},\quad V_{i}=\frac{1}{2\beta}\ln\left(\frac{k_{i}^{2}}{k_{i}^{2}-z_{i}^{2}}\right),\quad i=1,\dots,n$ (35)
where $k_{1}=k_{x1}-k_{xd}$, and $k_{i}$, for $i=2,\dots,n$, are positive constants, and $\beta>0$ is a design parameter. $V$ is positive definite and continuously differentiable in the set $|z_{i}|<k_{i}$ for all $i=1,2,\dots,n$.
The stabilizing functions and control law are designed as:
$\alpha_{1}=\frac{1}{g_{1}}\left(-f_{1}-\beta\kappa_{1}z_{1}\left(k_{1}^{2}-z_{1}^{2}\right)+\dot{y}_{d}\right)$ (36)
[…]
$u=\alpha_{n}$ (38)
This yields the closed-loop system:
$\dot{z}_{1}=-\beta\kappa_{1}z_{1}\left(k_{1}^{2}-z_{1}^{2}\right)+g_{1}z_{2}$ (39)
$\begin{split}\dot{z}_{i}&=-\beta\kappa_{i}z_{i}\left(k_{i}^{2}-z_{i}^{2}\right)-\frac{g_{i-1}z_{i-1}}{\beta}\frac{k_{i}^{2}-z_{i}^{2}}{k_{i-1}^{2}-z_{i-1}^{2}}\\ &+g_{i}z_{i+1},\quad i=2,\dots,n-1\end{split}$ (40)
$\dot{z}_{n}=-\beta\kappa_{n}z_{n}\left(k_{n}^{2}-z_{n}^{2}\right)-\frac{g_{n-1}z_{n-1}}{\beta}\frac{k_{n}^{2}-z_{n}^{2}}{k_{n-1}^{2}-z_{n-1}^{2}}$ (41)
The time derivative of $V$ can be rewritten as:
$\dot{V}=-\sum_{j=1}^{n}\kappa_{j}z_{j}^{2}$ (42)
Theorem 2: Consider the nonlinear system described by (39)–(41) with the control law (38) under Assumption 1, and assume the initial conditions satisfy $|z_{i}(0)|<k_{i}$ for $i=1,2,\dots,n$. Let the design parameters $\beta>0$ and $\kappa_{i}>0$ be chosen appropriately. Then, the closed-loop system ensures the following properties:
(i) The signals $z_{i}(t)$, $i=1,\dots,n$, remain bounded within the compact set $\Omega_{z}$, where $D_{z_{i}}$ is defined as:
$D_{z_{i}}=k_{i}\sqrt{1-e^{-2\beta V(0)}\prod_{j=1}^{n}\left(\frac{k_{j}^{2}-z_{j}^{2}(0)}{k_{j}^{2}}\right)}$
for all $i=1,\dots,n$.
(ii) The full states $x_{i}(t)$, $i=1,\dots,n$, remain within the constraint set $|x_{i}(t)|<k_{xi}$ for all $t\geq 0$.
(iii) All closed-loop signals are bounded.
(iv) The tracking error $z_{i}(t)=x_{i}(t)-x_{d}(t)$ converges to zero asymptotically as $t\to\infty$, i.e., $\lim_{t\to\infty}z_{1}(t)=0$, and the states $z_{i}(t)\to 0$, for $i=2,\dots,n$.
Proof:
i) Since $\dot{V}\leq 0$, we have $V(t)\leq V(0)$ for all $t\geq 0$, implying:
$\frac{1}{2\beta}\ln\left(\frac{k_{i}^{2}}{k_{i}^{2}-z_{i}^{2}}\right)\leq\sum_{j=1}^{n}\frac{1}{2\beta}\ln\left(\frac{k_{j}^{2}}{k_{j}^{2}-z_{j}^{2}(0)}\right)$
Thus, $|z_{i}(t)|\leq D_{z_{i}}$ holds for all $t\geq 0$, where $i=1,\dots,n$.
(ii) From the boundedness of $z_{1}(t)$ in (i), we have: $|x_{1}(t)|=|z_{1}(t)+x_{d}(t)|\leq D_{z_{1}}+|x_{d}(t)|$. Since $|x_{d}(t)|\leq k_{xd}$, it follows that: $|x_{1}(t)|\leq D_{z_{1}}+k_{xd}=k_{x1}$. For $i=2,\dots,n$, since $x_{i}(t)=z_{i}(t)+\alpha_{i-1}(t)$, the boundedness of $z_{i}(t)$ and $\alpha_{i-1}(t)$ (from (iii) below) ensures that the states $x_{i}(t)$ remain bounded within the constraint set $|x_{i}(t)|<k_{xi}$.
(iii) From (i), the error signals $z_{1}(t),\dots,z_{n}(t)$ are bounded. Considering Assumption 1, and the boundedness of $z_{1}(t)$ imply that $x_{1}(t)$ is also bounded. From (36), the stabilizing function $\alpha_{1}(t)$ is bounded, and thus, $x_{2}=z_{2}+\alpha_{1}$ is bounded. Similarly, from (37), the stabilizing functions $\alpha_{i}(t)$, for $i=2,\dots,n-1$, are bounded due to their dependence on bounded terms, ensuring that the states $x_{i}(t)$, $i=2,\dots,n$, are bounded. Since $u=\alpha_{n}$ from (38), and $\alpha_{n}(t)$ is bounded, the control input $u(t)$ is also bounded. Therefore, all closed-loop signals are bounded.
(iv) From (i) and (iii), the signals $z_{i}(t)$, $i=1,\dots,n$, are bounded, which implies that $V(t)$ is bounded. Given that $\dot{V}(t)\leq 0$ is non-increasing, it follows that $\ddot{V}(t)$ is bounded. By applying Barbalat’s Lemma, we conclude that $\dot{V}(t)\to 0$ as $t\to\infty$, leading to $z_{i}(t)\to 0$ for $i=1,\dots,n$, ensuring that the tracking error $z_{1}(t)\to 0$ asymptotically as $t\to\infty$.
IV Simulation
We consider the second-order nonlinear system:
$\dot{x}_{1}=\Theta_{1}x_{1}^{2}+x_{2},\quad\dot{x}_{2}=\Theta_{2}x_{1}x_{2}+\Theta_{3}x_{1}+(1+x_{1}^{2})u,$
where $\Theta_{1}=0.1$, $\Theta_{2}=0.1$, and $\Theta_{3}=-0.2$ [1]. The control design employs a s-ZBLF for output constraints, with $k_{1}=0.56$. The objective is to track the desired trajectory $y_{d}=0.2+0.3\sin(t)$, ensuring that the state $x_{1}$ remains within the safe constraint $|x_{1}|<k_{1}$. The initial conditions are set as $x_{1}(0)=0.25$ and $x_{2}(0)=1.5$. The control parameters $\kappa_{1}=2$ and $\kappa_{2}=2$ are used to adjust the response, while $\beta=10$ governs the control effort transition near the boundaries.
Figures 1 and 2 show the states $x_{1}$ and $x_{2}$ trajectories over time, with $x_{1}$ constrained by $|x_{1}|<0.56$. Figure 3 illustrates the control input $u$ applied to the system. Figure 4 compares the errors $z_{1}$ and $z_{2}$, while Figure 5 presents the relationship between $z_{1}$ and $z_{2}$. The results demonstrate that the control law ensures constraint satisfaction and smooth system behavior.
Figure 1: State $x_{1}$ over time with constraints $|x_{1}|<0.56$.
Figure 2: State $x_{2}$ over time.
Figure 3: Control input $u$ over time.
Figure 4: Comparison of errors $z_{1}$ and $z_{2}$ over time.
Figure 5: Relationship between errors $z_{1}$ and $z_{2}$.
V Conclusion
In this paper, we introduced the s-ZBLF for output and full-state constrained nonlinear control systems. The method offers a novel approach by maintaining minimal control effort near the origin and smoothly increasing it as the system state moves toward the constraint boundaries, providing a continuous and stable control response. This gradual transition avoids abrupt changes in control effort, addressing limitations found in earlier methods, and improves system performance by reducing chattering. Two specific forms—logarithmic-based and rational-based—were proposed, each demonstrating their effectiveness through theoretical analysis and simulation results for output-constrained systems. Future work will focus on optimizing the parameter $\beta$ to fine-tune the rate of control effort increase as the system state approaches the constraints. Additionally, we will explore the extension of s-ZBLF to multi-input multi-output (MIMO) systems and adaptive control schemes, as well as its application to systems with partial, asymmetric, time-varying, or unknown constraints to further enhance its practical applicability.

### v2

@@
-Smooth Zone Barrier Lyapunov Functions for Nonlinear Constrained Control Systems
+Progressive Barrier Lyapunov Functions for Nonlinear Constrained Control Systems
 Abstract
-This paper introduces the Smooth Zone Barrier Lyapunov Function (s-ZBLF) for output and full-state constrained nonlinear control systems.
-Unlike traditional BLF methods, where control effort continuously increases as the state moves toward the constraint boundaries, the s-ZBLF method keeps the control effort nearly zero near the origin, with a more aggressive increase as the system approaches the boundary.
-However, unlike previous works where control effort was zero within a predefined safe region around the origin, the s-ZBLF overcomes the disadvantage of discontinuous control activation by providing a smooth, gradual increase in control effort as the state nears the constraints.
-This smooth transition improves continuity in the control response and enhances stability by reducing chattering.
-Additionally, the s-ZBLF provides the advantage of minimal control effort in regions far from the constraints, reducing energy consumption and actuator wear.
-Two forms of the s-ZBLF—logarithmic-based and rational-based—are presented.
+This paper introduces the Progressive Barrier Lyapunov Function (p-BLF) for output- and full-state-constrained nonlinear control systems.
+Unlike traditional BLF methods, where control effort continuously increases as the state approaches the constraint boundaries, the p-BLF maintains minimal control effort in unconstrained regions and increases it progressively toward the boundaries.
+In contrast to previous methods with predefined safe zones and abrupt control activation, the p-BLF provides a smooth transition, improving continuity in the control response and enhancing stability by reducing chattering.
+Two forms of the p-BLF, logarithmic-based and rational-based, are developed to handle systems with either output constraints or full-state constraints.
 Theoretical analysis guarantees that all system states remain within the defined constraints, ensuring boundedness and stability of the closed-loop system.
-Simulation results validate the effectiveness of the proposed method in handling constrained nonlinear systems.
+Simulation results validate the effectiveness of the proposed method in constrained nonlinear control.
 Index Terms:
-Zone Barrier Lyapunov Function, Constrained Control, Nonlinear System, Adaptive Control, Full State Constraints.
+Progressive Barrier Lyapunov Function, Constrained Control, Nonlinear System, Adaptive Control, Full State Constraints.
 I Introduction
@@
 To address these constraints, Barrier Lyapunov Functions (BLFs) have emerged as a powerful tool, providing a systematic approach to enforce hard constraints on the system states [2, 3, 7].
-To overcome these limitations, the Zone Barrier Lyapunov Function (zBLF) was introduced, offering an improvement over traditional BLFs by introducing the concept of zones [4].
-In zBLF, system states are allowed to move freely within predefined safe zones, and control effort is only applied as the states approach the boundary of these zones.
-This significantly reduces unnecessary control actions, improves energy efficiency, and reduces actuator wear.
-The zBLF method has shown great promise, particularly in applications where maintaining efficient control is important, as it allows the system to operate without intervention within safe zones.
-Despite the advantages of zBLF, there are some inherent limitations.
-Most notably, zBLF suffers from discontinuous control activation near the boundary.
-As the system states transition from the free-moving zone to the constrained region, the control effort can switch on abruptly, leading to sharp control actions.
-This lack of fine-grained control near the boundaries can result in chattering, control instability, or degraded system performance, particularly in applications requiring smooth and continuous control.
-Additionally, zBLF’s reliance on strict boundaries can make it challenging to ensure a smooth and gradual increase in control effort as the system approaches the constraints.
-To address these challenges, we propose a new approach called the Smooth Zone Barrier Lyapunov Function (s-ZBLF).
-The s-ZBLF retains the energy-saving advantages of zBLF while addressing the issue of discontinuous control activation.
-Our approach introduces a smooth transition between the safe zone and the constrained region, providing a gradual increase in control effort as the system state moves closer to the boundary.
-By making the control effort grow progressively, s-ZBLF ensures continuity in the Lyapunov derivative, which leads to smoother and more stable control behavior.
-Near the origin, where the error is low, the control effort remains minimal, gradually increasing as the system approaches the constraint boundaries, and applying hard constraints only when the boundary is near.
-This progression ensures that control actions are smoother, leading to more effective constraint enforcement while maintaining system stability.
+Traditional BLF-based control strategies have been extensively studied for enforcing state constraints in nonlinear systems.
+Various forms of BLFs have been proposed, including tangent-type [9], integral-type [5], and logarithmic-type functions [1, 6], all designed to increase rapidly as the state approaches the constraint boundary.
+While this property ensures strong constraint enforcement, it also leads to high control effort even when the system states remain far from the boundaries.
+Such unnecessary control activation can increase energy consumption and accelerate actuator wear, particularly in applications where the constraints are rarely active.
+To overcome the high control effort associated with traditional BLFs, the Zone Barrier Lyapunov Function (zBLF) was introduced as an alternative that defines a predefined safe region around the origin [4, 8].
+Within this zone, system states are allowed to evolve freely without control activation, and corrective effort is applied only as the state approaches the zone boundary.
+This structure helps reduce unnecessary control actions, improve energy efficiency, and limit actuator usage.
+However, zBLF presents limitations related to its discontinuous control behavior at the zone boundary.
+The abrupt activation of control introduces nonsmoothness in the system’s response, which can lead to discontinuities in the velocity and sharp variations in acceleration.
+These effects may cause chattering, degraded tracking performance, and increased mechanical stress in applications that require smooth and continuous dynamics.
+Moreover, the fixed zone definition complicates the design of controllers that aim to provide a gradual and continuous increase in effort as the constraint boundary is approached.
+To address these challenges, we propose a new class of BLFs, referred to as the Progressive Barrier Lyapunov Function (p-BLF).
+The p-BLF method retains the low-effort behavior of zone-based approaches but removes the need for explicit zone definitions.
+It provides a gradual increase in control effort as the system state approaches the constraint boundaries, avoiding nonsmooth control activation.
+This progression ensures continuity in the Lyapunov derivative and improves stability by reducing chattering and discontinuities in the system response.
+The control effort remains minimal in unconstrained regions and increases only as needed to enforce the constraints.
 The key contributions of this paper are as follows:
-  • We propose the s-ZBLF, which introduces smooth transitions in control effort as the system approaches constraint boundaries, addressing the discontinuous control issue present in zBLF.
-  • s-ZBLF maintains minimal control effort when the system state is far from the constraint boundaries, while gradually increasing the control effort as the state moves closer to the constraints, ensuring smooth and stable transitions.
-  • The proposed method is demonstrated through theoretical analysis and simulation results, showing the performance of smooth control, stability, and constraint handling of s-ZBLF.
+  • A new class of BLFs, called p-BLF, is introduced to enable gradual and continuous control activation near constraint boundaries, addressing nonsmoothness and discontinuities in existing methods.
+  • The p-BLF framework maintains minimal control effort in unconstrained regions and increases it progressively as the state approaches the constraints, improving control efficiency and smoothness.
+  • Theoretical analysis and simulations demonstrate that the proposed approach ensures constraint satisfaction, closed-loop stability, and continuous control performance.
+The rest of the paper is structured as follows.
+Section II presents the problem formulation and relevant preliminaries.
+Section III introduces the proposed p-BLF, detailing its logarithmic-based and rational-based forms, and presents a control design example for a second-order SISO system.
+The section further extends the control design to higher-order systems with output and full-state constraints.
+Section IV presents simulation results, and Section V concludes the paper.
 II Problem Formulation and Preliminaries
 System Description: We consider a class of nonlinear systems in strict-feedback form:
-$\dot{x}_{i}=f_{i}(x_{1},\dots,x_{i})+g_{i}(x_{1},\dots,x_{i})x_{i+1},\quad i=1,\dots,n-1,$ (1)
-$\dot{x}_{n}=f_{n}(x_{1},\dots,x_{n})+g_{n}(x_{1},\dots,x_{n})u,$ (2)
+$\dot{x}_{i}$ $=f_{i}(x_{1},\dots,x_{i})+g_{i}(x_{1},\dots,x_{i})x_{i+1},$ $i=1,\dots,n-1,$
+$\dot{x}_{n}$ $=f_{n}(x_{1},\dots,x_{n})+g_{n}(x_{1},\dots,x_{n})u,$
+$y$ $=x_{1}.$ (1)
 where $x_{i}\in\mathbb{R}$ represents the system states, $u\in\mathbb{R}$ is the control input, and $y=x_{1}$ is the system output.
@@
 Then, $\lim_{t\to\infty}\dot{f}(t)=0$.
-Control Objective: The goal is to design a control law $u(t)$ such that the system tracks the desired trajectory $x_{d}(t)$, while ensuring the state constraints $|x_{i}(t)|<k_{xi}$ are satisfied, with $i=1$ for output constraints and $i=1,\dots,n$ for full-state constraints.
-Additionally, all closed-loop signals must remain bounded for $t\geq 0$.
+Control Objective: Given the nonlinear system described by (II), the objective is to design a control law $u(t)$ such that the output $y(t)=x_{1}(t)$ tracks a bounded desired trajectory $x_{d}(t)$, while satisfying the state constraints.
+Two scenarios are considered: (i) output constraint $|x_{1}(t)|<k_{x1}$, and (ii) full-state constraints $|x_{i}(t)|<k_{xi}$, $i=1,\dots,n$.
+To achieve this, a p-BLF is developed and integrated into the control design to ensure constraint satisfaction, closed-loop stability, and asymptotic tracking, while maintaining boundedness of all signals for $t\geq 0$.
 Barrier Lyapunov Functions [1] is a scalar function $V(x)$, defined with respect to the system $\dot{x}=f(x)$ on an open region $D$ containing the origin.
@@
 Along the solution of $\dot{x}=f(x)$, for $x(0)\in D$, it satisfies $V(x(t))\leq b$ for all $t\geq 0$, where $b$ is some positive constant.
-There are several types of BLF, such as tangent-type [5], integral-type [6], and logarithmic-type [7].
 One example of a standard BLF is given by the logarithmic form $V(x)=\frac{1}{2}\ln\left(\frac{k^{2}}{k^{2}-x^{2}}\right)$, where $k$ represents the constraint boundary.
@@
 This modification eliminates abrupt changes in control activation, providing smoother control transitions and enhancing system stability and performance.
-III Control Design with Smooth Zone Barrier Lyapunov Function
-III-A Smooth Zone Barrier Lyapunov Function
-Smooth Zone Barrier Lyapunov Function is introduced to provide smooth and continuous control activation, ensuring minimal control effort near the origin and progressively increasing control effort as the state nears the boundary.
-Unlike ZBLF, where control activation may be abrupt, s-ZBLF provides a gradual increase in control effort as the system state approaches the boundary, eliminating discontinuities and enhancing system stability.
-For a system state $x$ with constraint $|x|<k$, the following two examples of s-ZBLF are considered:
+III Control Design with Progressive Barrier Lyapunov Function
+III-A Progressive Barrier Lyapunov Function
+The Progressive Barrier Lyapunov Function (p-BLF) is introduced to enable smooth and continuous control effort modulation.
+It ensures minimal control activation near the origin and progressively increases the effort as the system state approaches the constraint boundary.
+In contrast to zone-based BLFs, which may cause abrupt control switching, the p-BLF provides a gradual transition, improving continuity and closed-loop stability.
+For a system state $x$ subject to the constraint $|x|<k$, two example forms of the p-BLF are considered:
 1. $V(x)=\frac{1}{2\beta}\ln\left(\frac{k^{2}}{k^{2}-x^{2}}\right)$
@@
 The objective is to design an output constrained control law $u$ that respects the state constraint $|x_{1}(t)|<k_{x1}$ while maintaining system stability.
-We illustrate both forms of the s-ZBLF presented in the previous section with examples.
-III-B1 Logarithmic-Based s-ZBLF
+We illustrate both forms of the p-BLF presented in the previous section with examples.
+III-B1 Logarithmic-Based p-BLF
 The logarithmic Lyapunov function is given by:
@@
 This ensures $\dot{V}_{2}\leq 0$, establishing the stability of the closed-loop system, and the boundedness of $x_{1}$ considering Assumption 1, and Lemma 2.
-III-B2 Rational-Based s-ZBLF
-For the rational-based s-ZBLF, the Lyapunov function is defined as:
+III-B2 Rational-Based p-BLF
+For the rational-based p-BLF, the Lyapunov function is defined as:
 $V_{1}=\frac{z_{1}^{2}}{2(k^{2}-z_{1}^{2})(1+\beta z_{1}^{2})}$ (15)
@@
 $\dot{V}_{2}=-\kappa_{1}z_{1}^{2}-\kappa_{2}z_{2}^{2}.$ (22)
-This ensures $\dot{V}_{2}\leq 0$, establishing the stability of the closed-loop system, and the boundedness of $x_{1}$ considering Assumption 1, and Lemma 2 for the rational-based s-ZBLF.
-III-C s-ZBLF for Higher-Order Nonlinear Systems with Output Constraints
+This ensures $\dot{V}_{2}\leq 0$, establishing the stability of the closed-loop system, and the boundedness of $x_{1}$ considering Assumption 1, and Lemma 2 for the rational-based p-BLF.
+III-C p-BLF for Higher-Order Nonlinear Systems with Output Constraints
 Backstepping is used in the control design in this section.
@@
 Define the state constraint as $|x_{1}(t)|<k_{x1}$, and let $k_{1}=k_{x1}-k_{xd}$ is the error constraint.
-The Lyapunov function candidates are defined using the logarithmic-based s-ZBLF as follows:
+The Lyapunov function candidates are defined using the logarithmic-based p-BLF as follows:
 $V_{1}=\frac{1}{2\beta}\ln\left(\frac{k_{1}^{2}}{k_{1}^{2}-z_{1}^{2}}\right)$ (23)
@@
 By applying Barbalat’s Lemma, we conclude that $\dot{V}_{n}(t)\to 0$ as $t\to\infty$, which leads to $z_{i}(t)\to 0$, for $i=1,\dots,n$, ensuring the output tracking error $z_{1}(t)$ converges to zero asymptotically.
-III-D s-ZBLF for Higher-Order Nonlinear Systems with Full State Constraints
+III-D p-BLF for Higher-Order Nonlinear Systems with Full State Constraints
 Denote $z_{1}=x_{1}-x_{d}$ and $z_{i}=x_{i}-\alpha_{i-1}$ for $i=2,\dots,n$. Consider the Lyapunov function candidate:
@@
 where $\Theta_{1}=0.1$, $\Theta_{2}=0.1$, and $\Theta_{3}=-0.2$ [1].
-The control design employs a s-ZBLF for output constraints, with $k_{1}=0.56$.
+The control design employs a p-BLF for output constraints, with $k_{1}=0.56$.
 The objective is to track the desired trajectory $y_{d}=0.2+0.3\sin(t)$, ensuring that the state $x_{1}$ remains within the safe constraint $|x_{1}|<k_{1}$.
@@
 V Conclusion
-In this paper, we introduced the s-ZBLF for output and full-state constrained nonlinear control systems.
+In this paper, we introduced the p-BLF for output and full-state constrained nonlinear control systems.
 The method offers a novel approach by maintaining minimal control effort near the origin and smoothly increasing it as the system state moves toward the constraint boundaries, providing a continuous and stable control response.
@@
 Future work will focus on optimizing the parameter $\beta$ to fine-tune the rate of control effort increase as the system state approaches the constraints.
-Additionally, we will explore the extension of s-ZBLF to multi-input multi-output (MIMO) systems and adaptive control schemes, as well as its application to systems with partial, asymmetric, time-varying, or unknown constraints to further enhance its practical applicability.
+Additionally, we will explore the extension of p-BLF to multi-input multi-output systems and adaptive control schemes, as well as its application to systems with partial, asymmetric, time-varying, or unknown constraints to further enhance its practical applicability.